A point charge q1=4.00 nC is placed at the origin, and a second point charge q2=-3.00 nC is

placed on the x-axis at x=+20.0 cm. A third point charge q3=2.00 nC is to be placed on the

x-axis between q1 and q2. Let the potential energy of the three charges be zero when they

are infinitely far apart.

a) what is the potential energy of the system of the three charges if q3 is placed at x=+10.0

cm?

b) where should q3 be placed to make the potential energy of the system equal to zero?


1
Expert's answer
2022-03-13T18:50:08-0400

a)

"E=E_1+E_2=kq_3(\frac{q_1}{x_1}+\frac{q_2}{x_2})=1.26~\mu J,"

b)

"\frac{q_1}{y}=\frac{q_2}{y-x},"

"y=80~cm."